The oscillations of two points x1 and x2 at x = 0 and x = 1 m respectively are modelled
as follows:
y1 = 0.2 sin 3πt
and y2 = 0.2 sin (3πt+ 8π)
Calculate the wavelength and speed of the associated wave.
1
Expert's answer
2019-02-28T11:40:08-0500

ω=3π rad/s.

kx=k(1)=8π

The wavelength:

λ=2π/k=2π/8π=0.25 m.

The speed of the associated wave

v=ω/k=3π/8π=0.75 m/s.